# In the pursuit for the nature of the $P_{c}(4457)$ and related pentaquarks

### v1

@@ Abstract
The magnetic moment of a hadron is an important spectroscopic parameter as its mass and encodes valuable information about its internal structure. In this present study, we systematically study magnetic moments of the $P_{c}(4457)$ and its related hidden-charm pentaquark states with and without strangeness employing a comprehensive analysis that encompasses diquark-diquark-antiquark scheme with $J^{P}=\frac{1}{2}^{-}$, $J^{P}=\frac{3}{2}^{-}$ and $J^{P}=\frac{5}{2}^{-}$ quantum numbers. Some of the obtained magnetic moments agree well with the available theoretical results. The predicted magnetic moment values together with alongside results existing in the literature, may offer insights into their underlying structures, and consequently their spin-parity quantum numbers.
@@ I Introduction
    Diquark-diquark-antiquark scheme: In this case, diquark prefers to form color antitriplet $\bar{3}_{c}$. Two color antriplet diquarks together with color antitriplet quark $\bar{3}_{c}(\text{diquark)}\otimes\bar{3}_{c}(\text{diquark)}\otimes\bar{3}_{c}(\text{antiquark)}$) form color singlet pentaquark.
@@
    Diquark-triquark scheme: This scheme has a similar configuration with respect to molecular scheme. In this type of configuration, triquark consists two quarks and antiquark. Color triplet triquark $3_{c}(\text{triquark)}$) and color antitriplet diquark $\bar{3}_{c}(\text{diquark)}$) form diquark-triquark configuration $3_{c}(\text{triquark)}\otimes\bar{3}_{c}(\text{diquark)}$) as color singlet a pentaquark state.
@@
Despite the fact that magnetic moment gives valuable information on physical parameters that are directly correlated with the inner structure of the state, the number of studies related to magnetic moment are not at the desired level [13, 14, 9, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Although the magnetic moments of the hadrons give valuable information about the internal structure, magnetic moment measurement is a real challenge. Magnetic moment is the response of a state to an external magnetic field and if the particle lifetime is relatively short, it is hard to measure. However, magnetic moment measurement is possible indirectly, provided sufficient data is accumulated. For example, $\Delta^{+}(1232)$ baryon has also a very short lifetime, $(\sim 5\times 10^{-24}\ \text{s})$, however its magnetic moment was obtained through $\gamma N$rightarrow $\Delta$rightarrow $\Delta\gamma$rightarrow $\pi N\gamma$ process [32, 33, 34]. A comparable process may be employed to achieve the magnetic moment of the $P_{c}$ pentaquarks. QCD based models can also be employed to obtain magnetic moments. In Refs. [35, 36], magnetic moments of baryons containing two charm quarks have been extracted by lattice QCD formalism. These analyses can be extended to the exotic hadrons.
@@ II.1 Wave function
[…]
@@
$(\frac{1}{2},\frac{1}{2})$ $-\sqrt{\frac{1}{3}}\{ud\}(cu){\bar{c}}+\sqrt{\frac{2}{3}}\{uu\}{({c}d)\bar{c}}$ $[ud]({c}u)\bar{c}$
$(\frac{1}{2},-\frac{1}{2})$ $\sqrt{\frac{1}{3}}\{ud\}(cd){\bar{c}}-\sqrt{\frac{2}{3}}\{dd\}({c}u){\bar{c}}$ $[ud](cd){\bar{c}}$
$(1,1)$ $\sqrt{\frac{1}{3}}\{us\}({c}u){\bar{c}}-\sqrt{\frac{2}{3}}\{uu\}(cs){\bar{c}}$ $[us](cu){\bar{c}}$
@@
$(1,-1)$ $\sqrt{\frac{1}{3}}\{ds\}({c}d){\bar{c}}-\sqrt{\frac{2}{3}}\{dd\}(cs){\bar{c}}$ $[ds](cd){\bar{c}}$
@@
$(\frac{1}{2},\frac{1}{2})$ $\sqrt{\frac{1}{3}}\{us\}({c}s){\bar{c}}-\sqrt{\frac{2}{3}}\{ss\}({c}u){\bar{c}}$ $[us](cs){\bar{c}}$
$(\frac{1}{2},-\frac{1}{2})$ $\sqrt{\frac{1}{3}}\{ds\}({c}s){\bar{c}}-\sqrt{\frac{2}{3}}\{ss\}({c}d){\bar{c}}$ $[ds](cs){\bar{c}}$
@@
$(1,-1)$ $\sqrt{\frac{2}{3}}\{ds\}(cd){\bar{c}}+\sqrt{\frac{1}{3}}\{dd\}(cs){\bar{c}}$
@@
where $S_{12}=1$ for symmetric flavor wave function and $S_{12}=0$ for antisymmetic wave function for the first diquark. This pattern follows for the second diquark, where $S_{34}=1$ or 00 for symmetric or antisymmetric flavor wave functions, respectively. $\ell$ represents the orbital angular momentum.
@@ III Numerical results and Discussion
$P_{c}(4457)$ $-\sqrt{\frac{1}{3}}\{ud\}(cu){\bar{c}}+\sqrt{\frac{2}{3}}\{uu\}{({c}d)\bar{c}}$ $[ud]({c}u)\bar{c}$
$P_{cr1}$ $\sqrt{\frac{1}{3}}\{ud\}(cd){\bar{c}}-\sqrt{\frac{2}{3}}\{dd\}({c}u){\bar{c}}$ $[ud](cd){\bar{c}}$
$P_{cr2}$ $\sqrt{\frac{1}{3}}\{us\}({c}u){\bar{c}}-\sqrt{\frac{2}{3}}\{uu\}(cs){\bar{c}}$ $[us](cu){\bar{c}}$
$P_{cr3}$ $\sqrt{\frac{1}{3}}\{ds\}({c}d){\bar{c}}-\sqrt{\frac{2}{3}}\{dd\}(cs){\bar{c}}$ $[ds](cd){\bar{c}}$
$P_{cr4}$ $\sqrt{\frac{1}{3}}\{us\}({c}s){\bar{c}}-\sqrt{\frac{2}{3}}\{ss\}({c}u){\bar{c}}$ $[us](cs){\bar{c}}$
$P_{cr5}$ $\sqrt{\frac{1}{3}}\{ds\}({c}s){\bar{c}}-\sqrt{\frac{2}{3}}\{ss\}({c}d){\bar{c}}$ $[ds](cs){\bar{c}}$
@@
$\frac{1}{2}^{-}$ ${{}^{2}S_{\frac{1}{2}}}$ ${0}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$0.377
${1}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$0.244
$\frac{3}{2}^{-}$ ${{}^{4}S_{\frac{3}{2}}}$ ${1}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$0.930
@@
$(1^{+}\otimes 1^{+})_{0}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$0.377
@@
$\frac{1}{2}^{-}$ ${{}^{2}S_{\frac{1}{2}}}$ ${0}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$0.377
@@
$\frac{1}{2}^{-}$ ${{}^{2}S_{\frac{1}{2}}}$ $0^{+}\otimes 1^{+}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$1.115
$(1^{+}\otimes 1^{+})_{0}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$0.377
@@
$\frac{3}{2}^{-}$ ${{}^{4}S_{\frac{3}{2}}}$ $(0^{+}\otimes 1^{+})\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$2.238
$(1^{+}\otimes 1^{+})_{1}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$0.188
@@
$\frac{1}{2}^{-}$ ${{}^{2}S_{\frac{1}{2}}}$ ${0}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$0.377
${1}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$0.009
$\frac{3}{2}^{-}$ ${{}^{4}S_{\frac{3}{2}}}$ ${1}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$0.579
@@
$(1^{+}\otimes 1^{+})_{0}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$0.377
@@
$\frac{1}{2}^{-}$ ${{}^{2}S_{\frac{1}{2}}}$ ${0}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$0.377
${1}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$0.009
$\frac{3}{2}^{-}$ ${{}^{4}S_{\frac{3}{2}}}$ ${1}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$0.579
@@
$\frac{1}{2}^{-}$ ${{}^{2}S_{\frac{1}{2}}}$ $0^{+}\otimes 1^{+}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$1.115
$(1^{+}\otimes 1^{+})_{0}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$0.377
$(1^{+}\otimes 1^{+})_{1}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$0.562
$\frac{3}{2}^{-}$ ${{}^{4}S_{\frac{3}{2}}}$ $(0^{+}\otimes 1^{+})\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$2.238
$(1^{+}\otimes 1^{+})_{1}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$1.408
$(1^{+}\otimes 1^{+})_{2}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$1.631
$\frac{5}{2}^{-}$ ${{}^{6}S_{\frac{5}{2}}}$ $1^{+}\otimes 1^{+}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$2.440
@@
$\frac{1}{2}^{-}$ ${{}^{2}S_{\frac{1}{2}}}$ ${0}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$0.377
@@
$\frac{1}{2}^{-}$ ${{}^{2}S_{\frac{1}{2}}}$ $0^{+}\otimes 1^{+}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$0.646
$(1^{+}\otimes 1^{+})_{0}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$0.377
@@
$\frac{3}{2}^{-}$ ${{}^{4}S_{\frac{3}{2}}}$ $(0^{+}\otimes 1^{+})\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$1.535
@@
$\frac{1}{2}^{-}$ ${{}^{2}S_{\frac{1}{2}}}$ ${0}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$0.377
${1}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$0.244
$\frac{3}{2}^{-}$ ${{}^{4}S_{\frac{3}{2}}}$ ${1}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$0.930
@@
$\frac{1}{2}^{-}$ ${{}^{2}S_{\frac{1}{2}}}$ $0^{+}\otimes 1^{+}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$0.646
$(1^{+}\otimes 1^{+})_{0}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$0.377
$(1^{+}\otimes 1^{+})_{1}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$0.445
$\frac{3}{2}^{-}$ ${{}^{4}S_{\frac{3}{2}}}$ $(0^{+}\otimes 1^{+})\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$1.535
$(1^{+}\otimes 1^{+})_{1}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$1.233
$(1^{+}\otimes 1^{+})_{2}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$1.315
$\frac{5}{2}^{-}$ ${{}^{6}S_{\frac{5}{2}}}$ $1^{+}\otimes 1^{+}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$2.088

### v2

# Magnetic Moments of Hidden-Charm Pentaquarks in the Diquark-Diquark-Antiquark Scheme

@@ Abstract
The magnetic moment of a hadron is an important spectroscopic parameter that encodes valuable information about its internal structure. In this work, we systematically investigate the magnetic moments of hidden-charm pentaquark states, including the experimentally observed $P_{c}(4457)$ and related configurations with and without strangeness. The analysis is performed within the diquark-diquark-antiquark framework for spin-parity quantum numbers $J^{P}=\frac{1}{2}^{-}$, $\frac{3}{2}^{-}$, and $\frac{5}{2}^{-}$. Magnetic moment values are computed for different spin and flavor configurations, and the results are compared with existing predictions in the literature. These predictions may offer insight into the inner structure and quantum numbers of these exotic states, and potentially help distinguish between different theoretical models.
@@ I Introduction
    Diquark-diquark-antiquark scheme: In this case, diquark prefers to form color antitriplet $\bar{3}_{c}$. Two color antriplet diquarks together with color antitriplet quark $\bar{3}_{c}(\text{diquark)}\otimes\bar{3}_{c}(\text{diquark)}\otimes\bar{3}_{c}(\text{antiquark)}$ form color singlet pentaquark.
@@
    Diquark-triquark scheme: This scheme has a similar configuration with respect to molecular scheme. In this type of configuration, triquark consists two quarks and antiquark. Color triplet triquark $3_{c}(\text{triquark)}$ and color antitriplet diquark $\bar{3}_{c}(\text{diquark)}$ form diquark-triquark configuration $3_{c}(\text{triquark)}\otimes\bar{3}_{c}(\text{diquark)}$ as color singlet a pentaquark state.
@@
Despite the fact that magnetic moment gives valuable information on physical parameters that are directly correlated with the inner structure of the state, the number of studies related to magnetic moment are not at the desired level [13, 14, 9, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Although the magnetic moments of the hadrons give valuable information about the internal structure, magnetic moment measurement is a real challenge. Magnetic moment is the response of a state to an external magnetic field and if the particle lifetime is relatively short, it is hard to measure. However, magnetic moment measurement is possible indirectly, provided sufficient data is accumulated. For example, $\Delta^{+}(1232)$ baryon has also a very short lifetime, $(\sim 5\times 10^{-24}\ \text{s})$, however its magnetic moment was obtained through $\gamma N$ $\rightarrow$ $\Delta$ $\rightarrow$ $\Delta\gamma$ $\rightarrow$ $\pi N\gamma$ process [32, 33, 34]. A comparable process may be employed to achieve the magnetic moment of the $P_{c}$ pentaquarks. QCD based models can also be employed to obtain magnetic moments. In Refs. [35, 36], magnetic moments of baryons containing two charm quarks have been extracted by lattice QCD formalism. These analyses can be extended to the exotic hadrons.
@@ II.1 Wave function
As noted in Eq. (3), diquarks can form symmetric $(6_{c})$ or antisymmetric $(\bar{3}_{c})$ color configurations. These configurations are allowed by color SU(3). The antisymmetric $(\bar{3}_{c})$ is chosen because it is the dominant configuration in most phenomenological models of compact pentaquarks, as it allows color-singlet formation with the antiquark. It is also energetically favored due to the attractive nature of the color interaction in this channel which is a well-established result in QCD-based models and supported by lattice QCD simulations and phenomenological studies (see Refs. [10, 11]). The symmetric $(6_{c})$ diquark would require additional color-neutral binding mechanisms (e.g., gluon exchanges), complicating the model without clear experimental justification.
@@
Here $q$ represents the light quarks, $u$ and $d$. In the above flavor wave functions, the bracket $\left\{~\right\}$" for the first diquark denotes that two quarks are symmetric (spin 1) inside it and the bracket $\left[~\right]$" for the first diquark denotes two quarks are antisymmetric (spin 0) inside it in the SU(3) flavor space. For the second diquark, the bracket ()``(~)" means that flavor symmetry of the two quarks inside this diquark is not definite. It is possible to use SU(2) Clebcsh-Gordon coefficients to expand the flavor wave functions in Eqs. (4-15) to obtain flavor wave functions in $8_{f}$ representation (Table 2) and $10_{f}$ representation (Table 3):
@@
$(\frac{1}{2},\frac{1}{2})$ $\sqrt{\frac{1}{3}}\{ud\}(cu){\bar{c}}+\sqrt{\frac{2}{3}}\{uu\}{({c}d)\bar{c}}$ $ud]({c}u)\bar{c}$
$(\frac{1}{2},-\frac{1}{2})$ $\sqrt{\frac{1}{3}}\{ud\}(cd){\bar{c}}-\sqrt{\frac{2}{3}}\{dd\}({c}u){\bar{c}}$ $ud](cd){\bar{c}}$
$(1,1)$ $\sqrt{\frac{1}{3}}\{us\}({c}u){\bar{c}}-\sqrt{\frac{2}{3}}\{uu\}(cs){\bar{c}}$ $us](cu){\bar{c}}$
@@
(1,1)(1,-1) $\sqrt{\frac{1}{3}}\{ds\}({c}d){\bar{c}}-\sqrt{\frac{2}{3}}\{dd\}(cs){\bar{c}}$ $ds](cd){\bar{c}}$
@@
$(\frac{1}{2},\frac{1}{2})$ $\sqrt{\frac{1}{3}}\{us\}({c}s){\bar{c}}-\sqrt{\frac{2}{3}}\{ss\}({c}u){\bar{c}}$ $us](cs){\bar{c}}$
$(\frac{1}{2},-\frac{1}{2})$ $\sqrt{\frac{1}{3}}\{ds\}({c}s){\bar{c}}-\sqrt{\frac{2}{3}}\{ss\}({c}d){\bar{c}}$ $ds](cs){\bar{c}}$
@@
(1,1)(1,-1) $\sqrt{\frac{2}{3}}\{ds\}(cd){\bar{c}}+\sqrt{\frac{1}{3}}\{dd\}(cs){\bar{c}}$
@@
where $S_{12}=1$ for symmetric flavor wave function and $S_{12}=0$ for antisymmetic wave function for the first diquark. This pattern follows for the second diquark, where $S_{34}=1$ or 0 for symmetric or antisymmetric flavor wave functions, respectively. $\ell$ represents the orbital angular momentum. In this work, we assume $\ell=0$, corresponding to S-wave pentaquark states, where the total orbital angular momentum includes the relative motion between the subclusters (diquark-diquark-antiquark).
@@ III Numerical results and Discussion
$P_{c}(4457)$ $\sqrt{\frac{1}{3}}\{ud\}(cu){\bar{c}}+\sqrt{\frac{2}{3}}\{uu\}{({c}d)\bar{c}}$ $ud]({c}u)\bar{c}$
$P_{cr1}$ $\sqrt{\frac{1}{3}}\{ud\}(cd){\bar{c}}-\sqrt{\frac{2}{3}}\{dd\}({c}u){\bar{c}}$ $ud](cd){\bar{c}}$
$P_{cr2}$ $\sqrt{\frac{1}{3}}\{us\}({c}u){\bar{c}}-\sqrt{\frac{2}{3}}\{uu\}(cs){\bar{c}}$ $us](cu){\bar{c}}$
$P_{cr3}$ $\sqrt{\frac{1}{3}}\{ds\}({c}d){\bar{c}}-\sqrt{\frac{2}{3}}\{dd\}(cs){\bar{c}}$ $ds](cd){\bar{c}}$
$P_{cr4}$ $\sqrt{\frac{1}{3}}\{us\}({c}s){\bar{c}}-\sqrt{\frac{2}{3}}\{ss\}({c}u){\bar{c}}$ $us](cs){\bar{c}}$
$P_{cr5}$ $\sqrt{\frac{1}{3}}\{ds\}({c}s){\bar{c}}-\sqrt{\frac{2}{3}}\{ss\}({c}d){\bar{c}}$ $ds](cs){\bar{c}}$
@@
$\frac{1}{2}^{-}$ $^{2}S_{\frac{1}{2}}$} ${0}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$ -0.377
${1}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$ -0.244
$\frac{3}{2}^{-}$ $^{4}S_{\frac{3}{2}}$} ${1}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$ -0.930
@@
$(1^{+}\otimes 1^{+})_{0}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -0.377
@@
$\frac{1}{2}^{-}$ $^{2}S_{\frac{1}{2}}$} ${0}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$ -0.377
@@
$\frac{1}{2}^{-}$ $^{2}S_{\frac{1}{2}}$} $0^{+}\otimes 1^{+}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -1.115
$(1^{+}\otimes 1^{+})_{0}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -0.377
@@
$\frac{3}{2}^{-}$ $^{4}S_{\frac{3}{2}}$} $(0^{+}\otimes 1^{+})\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -2.238
$(1^{+}\otimes 1^{+})_{1}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -0.188
@@
$\frac{1}{2}^{-}$ $^{2}S_{\frac{1}{2}}$} ${0}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$ -0.377
${1}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$ -0.009
$\frac{3}{2}^{-}$ $^{4}S_{\frac{3}{2}}$} ${1}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$ -0.579
@@
$(1^{+}\otimes 1^{+})_{0}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -0.377
@@
$\frac{1}{2}^{-}$ $^{2}S_{\frac{1}{2}}$} ${0}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$ -0.377
${1}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$ -0.009
$\frac{3}{2}^{-}$ $^{4}S_{\frac{3}{2}}$} ${1}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$ -0.579
@@
$\frac{1}{2}^{-}$ $^{2}S_{\frac{1}{2}}$} $0^{+}\otimes 1^{+}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -1.115
$(1^{+}\otimes 1^{+})_{0}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -0.377
$(1^{+}\otimes 1^{+})_{1}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -0.562
$\frac{3}{2}^{-}$ $^{4}S_{\frac{3}{2}}$} $(0^{+}\otimes 1^{+})\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -2.238
$(1^{+}\otimes 1^{+})_{1}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -1.408
$(1^{+}\otimes 1^{+})_{2}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -1.631
$\frac{5}{2}^{-}$ $^{6}S_{\frac{5}{2}}$} $1^{+}\otimes 1^{+}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -2.440
@@
$\frac{1}{2}^{-}$ $^{2}S_{\frac{1}{2}}$} ${0}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$ -0.377
@@
$\frac{1}{2}^{-}$ $^{2}S_{\frac{1}{2}}$} $0^{+}\otimes 1^{+}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -0.646
$(1^{+}\otimes 1^{+})_{0}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -0.377
@@
$\frac{3}{2}^{-}$ $^{4}S_{\frac{3}{2}}$} $(0^{+}\otimes 1^{+})\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -1.535
@@
$\frac{1}{2}^{-}$ $^{2}S_{\frac{1}{2}}$} ${0}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$ -0.377
${1}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$ -0.244
$\frac{3}{2}^{-}$ $^{4}S_{\frac{3}{2}}$} ${1}^{+}\otimes 0^{+}\otimes{\frac{1}{2}}^{-}\otimes{0^{+}}$ -0.930
@@
$\frac{1}{2}^{-}$ $^{2}S_{\frac{1}{2}}$} $0^{+}\otimes 1^{+}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -0.646
$(1^{+}\otimes 1^{+})_{0}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -0.377
$(1^{+}\otimes 1^{+})_{1}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -0.445
$\frac{3}{2}^{-}$ $^{4}S_{\frac{3}{2}}$} $(0^{+}\otimes 1^{+})\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -1.535
$(1^{+}\otimes 1^{+})_{1}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -1.233
$(1^{+}\otimes 1^{+})_{2}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -1.315
$\frac{5}{2}^{-}$ $^{6}S_{\frac{5}{2}}$} $1^{+}\otimes 1^{+}\otimes{\frac{1}{2}}^{-}\otimes 0^{+}$ -2.088
@@
  • The $J^{P}$ assignments in Tables 510 follow the conventions: the anti-charm quark has $J^{P}=1/2^{-}$ due to its fermionic nature and negative intrinsic parity; the orbital part is assumed to be in an S-wave ($J^{P}=0^{+}$) throughout this work.